a 10% efficient engine accelerates a 1500 kg car from rest to 15 m/s. how much energy is transferred to the engine by burning gasoline?
Answer #1

Data Given

mass of the car m = 1500 kg, initial velocity vi = 0 m/s, final velocity vf = 15 m/s

efficiency n = 10%

Now output is equal to work done and we will calculate work done using work energy theorem.

According to work energy theorem work done = change in kinetic energy


W = \frac{1}{2}mv_{f}^{2}-\frac{1}{2}mv_{i}^{2}

W = = x 1500kg < (15m/s)2 – 3 x 1500kg < (Om/s)


W = 168750J

Now we know that efficiency is given by


Output - X 100 Input

So

168750J 10 x 100 = 100 X 100 Einput


168750J Einput 10 100


100 Einput = 168750J X = 1687500) 10

Einput = 1687.5 kJ